11-065 High Voltage Arcing problem and solution

Print Friendly, PDF & Email

As the 11-065s get older we are starting to see instances where the beam voltage, condenser voltage or objective voltage becomes unstable as the front panel potentiometers are adjusted.

The front panel potentiometers can become “noisy” as a result of oxidation on the internal contacts.  In addition to causing instability in the output voltages, it is also possible for the potentiometer output potentiometer “open” up.  When this happens, the output of the beam, condenser or objective supply can go as high as 6.4kV as shown in the picture below in which a 1000:1 high voltage probe is used to measure the beam voltage supply.

In the case of the condenser or objective supplies becoming unstable the result is that the ion beam might go in and out of focus, or the ion beam can get completely shut off.  However, if the beam voltage becomes unstable and goes up to 6.5kV then the opto-couplers on the HV1 board will become damaged and then the emission or pressure circuits will no longer function properly.  The opto-couplers are only rated up to 5kV.

If you suspect that your 11-065 beam, condenser or objective supplies are unstable, here is how to test the HV1 board outputs;

  1. Turn off the 11-065 and unplug the input power cord.
  2. Place the 11-065 on the bench and remove the top cover
  3. Remove the HV cover (on the right hand side of the unit)
  4. Unplug all of the spade connectors which connect the various wires to the HV1 and HV2 boardsHV1 outputs
  5. Lift out the HV2 board (the one closer to the center of the unit)
  6. Place all of the wires off to the side of the 11-065, making sure that none of them touch the chassis. Most of the wires are outputs and so have no voltage on them, but the 4 center wires on the HV2 board have 20VAC on them.
  7. Connect a high voltage probe to the beam output connector on the still plugged in HV1 board, (the board closest to the chassis) ground reference is the chassis. The outputs are:   E93 Beam Voltage, E90 OBJ, E89 COND.Beam voltage HV1 board
  8. Plug in the 11-065 input power cord.
  9. Make sure that the Beam voltage switch is OFF and the beam voltage knob is turned fully CCW.
  10. Turn on the 11-065 power.
  11. Turn on the beam voltage and monitor the output on the DVM that is connected to the high voltage probe.
  12. Slowly turn the Beam voltage potentiometer CW and observe the DVM reading. The Beam voltage output should increase smoothly from near zero to 5kV as you turn up the potentiometer.  If you see jumping, instability, or if the Beam voltage output goes up to 6.5 kV then the potentiometer is noisy and needs to be replaced.  The potentiometer is a 10 k ohm 5 turn 2 Watt 1% potentiometer available from DIgikey, Mouser and Newark.  
  13. The OBJ and COND outputs go from 50% to 100% of the Beam Voltage.  So to test those, the Beam voltage needs to stay fully CW at 5.0   The OBJ and COND potentiometers are also 10 k ohm 5 turn 2 watt 1% potentiometers. 

If the Beam voltage potentiometer was noisy and the voltage went higher than 5kV, then the opto-couplers were likely damaged.   If your do not get any emission current, then most likely U6 on the HV2 board was damaged and should be replaced.  Other components may be damaged as well. If the emission works but not the pressure, U7 is likely damaged.

On the really old 11-065s, U6 was a TIL109 opto.  When that part became obsolete about 20 years ago, it was replaced with a SPX314 opto.  Most of the 11-065s in the field have been updated to the SPX314 (a modification is required).   If you have a really old 11-065 with the TIL109 opto-couplers, RBD still has a few of those in stock.

Recently, the SPX314 has become obsolete and is hard to find.  It can be replaced with a OPI 110 opto-coupler which is also has 15kV of isolation, so even if the HV1 board goes up to 6.4kV this opto-coupler will not get damaged. The OPI 110 will not only repair the problem, it will ensure that this particular problem does not happen again.

There is no modification required to switch from a SPX314 to an OPI 110, but the pin outs are slightly different and are shown in the pictures below.

SPX314 on HV2 board
11-065 opto
OPI 110 opto connections
OPI 110 connection info
U6 schematic HV2 board

The OPI 110 opto-couplers are available from most major electronic part suppliers including Mouser.

Here is a link to the datasheet – OPI 110 datasheet

If you need technical help to repair your 11-065 ion gun control you can contact RBD Instruments Inc. for assistance. We offer technical support, repair/ calibration and the use of a loaner 11-065 while your unit is being repaired.

11-010 Electron Gun Control modification to reduce noise

Print Friendly, PDF & Email

Due to aging components, many of the PHI 11-010 5kV electron gun controls have developed a slight noise problem in the emission chopper circuit on the 623 board that results in about 300 mV of 60 hz ripple on the emission voltage, which translates into noise in the electron beam and Auger data.

Since the chopper circuit is never used (it was designed as a way to get N/E data with a lock-in recorder), a very simple solution to the problem is to bypass the chopper circuit.

Procedure:

  1. Unplug the 11-010 AC power cord and remove the cover.
  2. Solder a jumper between Pins 17 and 18 on the 623 board as shown below.  The 623 board is located on the side of the 11-010, just above the filament supply board.

This modification will reduce the noise level from about 300 mV to 50 mv or less and results in noticeably cleaner Auger data.

Chopper board resistors
This jumper will bypass the Chopper board circuitry
Chopper board schematic
Jumper location on Chopper board

20-610 3kV Adjustment

Print Friendly, PDF & Email

The 20-610 high voltage gun supply used on PHI 600 and 660 scanning Auger systems provides the beam voltage, filament current and emission voltage to the Lab6 filament in the cylindrical mirror analyzer.

As part of the linearity adjustment process, the beam voltage is set to 3 keV so that the top of the elastic peak comes in at 3 keV.

The 3 keV elastic peak adjustment potentiometer in the 20-610 is R108.   The process is simple:

  1. Adjust the specimen stage Z axis for maximum counts and best shape of the peak during a 3 keV elastic peak alignment.
  2. Adjust the analyzer control gain so that the AES peaks come in at the correct location in a survey. Typically clean copper is used since it has both low and high energy peaks.
  3. Reacquire the elastic peak but do not move the specimen stage. If the elastic peak does not come it at 3 keV, move the elastic peak to 3 keV by adjusting R108 in the 20-610.

Sometimes there is not enough range of adjustment with the R108 potentiometer.  For those cases, this blog post will explain how to modify the PCB 100 board in the 20-610 in order to extend the 3 kV adjustment range.

The schematic below shows the R108 potentiometer circuit.

R108 Pot

The ends of R108 are connected to + and – 12V via two 49.9 K ohm resistors, R107 and R109. Adjusting R108 adds a small offset current to IC 103, which in turn changes the 3 kV output of the 20-610.

If there is not enough range with R108 then we need to change the balance between the + and – 12 volt supplies.   The way to do that is to remove R107 and replace it with a 100 k ohm trim potentiometer that is set to an initial value of 49.9 K ohms. 100 K ohm trim pots are available from any electronic component vendor (Digikey, Newark, Mouser….)

To prep the 100 K ohm trim potentiometer, the center wiper connection needs to be soldered to either end as shown below.

Bend center pin

With an ohmmeter, measure the resistance between the trim pot legs and adjust it for a resistance of 49.9 K ohms.

Turn off the 20-610 and on the back of the unit, unplug the main power cord, the remote program cable and the unscrew and remove the large HV cable connector. If there is a ground wire attached, remove it as well.

Place the 20-610 on a work bench and remove the 100 board. You will need to remove the board tie down bracket. Important! – Make sure that the 20-610 in unplugged when removing or installing the board tie down bracket as it is very close to some exposed wires that have voltage on them when the 20-610 is plugged in.

 

R100 board location in 20-610

remove bracket

 

 

 

Unsolder R107 and install the new trim pot on the back side of the 100 board.   You will need to bend the two pins on the trim pot so that the adjustment screw faces up.

R107 resistor

Reinstall the 100 board inside the 20-610.  Make sure that it is seated all the way into the mother board connector. Reinstall the board tie down bracket. Do not put the cover back on the 20-610.

Install the 20-610 into the electronics rack and reconnect the large HV cable (be sure to screw it in all the way), and the HV programming cable.  Make sure that the 20-610 main power switch is  OFF and then plug in the main power cord into the interlocked power strip.

Slide the 20-610 out enough so that you can get to R108 and the new trim pot.  Make sure that the 20-610 main power is OFF.

Remove the filament cap from the analyzer (3 screws) and connect a DVM and high voltage probe to either of the filament tabs.  The ground reference on the high voltage probe should be connected to the vacuum chamber.

 

***Caution, high voltage is present! Refer adjustment to qualified personnel ***

Turn on the 20-610 and using AugerMap and the normal turn on procedure, set the beam voltage to 3 kV.

***Caution, high voltage is present! Refer adjustment to qualified personnel ***

Center R108 and then adjust the new 100 K ohm trim pot so that the voltage on the filament cap is 3,000 volts.

Turn off the beam voltage in AugerMap and then turn off the 20-610.

Reconnect the filament cap to the analyzer.  Press down on the cap as you tighten the 3 screws that hold the cap onto the top of the CMA filament adjustment housing.

Turn the 20-610 back on and turn the beam voltage on and set it to 3 kV.  Bring up the filament current to the normal operating point (typically 1.3 amps)

Perform the AES calibration procedure and adjust R108 so that the elastic peak comes in at 3 keV.   When you are finished with the AES calibration, put the cover back on the 20-610 (front 2 cover screws only) and slide it back into the console.  The AES calibration procedure is listed below –

Auger energy calibration on 600 and 660 scanning Auger systems

This procedure requires sliding the 20-610 high voltage supply out and removing the cover to gain access to the beam voltage offset potentiometer, R108. Turn off the 20-610 when sliding it in out or in, and when removing or installing the cover.

Procedure:

  1. Load a sample of pure copper.
  2. If you are using AugerMap software, set the magnification to 10,000X and use the Area Scan mode to minimize sample topography effect on the Auger signal.
  3. Perform an elastic peak alignment and adjust the Z axis sample position to obtain maximum counts and best peak shape.
  4. Sputter the sample clean until no carbon or oxygen is present.
  5. Re-acquire the elastic peak to ensure that the sample is at the optimum position: highest counts and best peak shape. When the elastic peak is differentiated, the positive and negative excursions should be equal and symmetrical.
  6. From this point on, do not move the sample!
  7. With the beam voltage at 3kV, acquire a survey from 30eV to 1030eV, using .5eV/step, 50 ms/point.
  8. Differentiate the survey and check the peak positions against the correct values as listed in the PHI handbook or other reference. A typical value is 920eV for the high energy peak and 60eV for the low energy peak on copper.
  9. Note: If using AugerScan software, you can simply adjust the scale factor in the AES
  10. Acquire an alignment with a range of 900 to 940, .5eV/step, 15ms/point and do the adjustment in real time. For copper, set the n/e peak to approximately 917eV. When differentiated, the high energy Cu peak should be 920eV.
  11. Acquire another survey and check that the differentiated peak positions are correct. Document the results for future reference and file it in the system calibration log.
  12. Acquire another elastic peak, but do not move the sample!
  13. If the elastic peak is not centered at 3kV, then adjust R108 in the Bertan 20-610 High Voltage power supply to center the elastic peak.

Calibration is complete.

From this point on, every-time you set the elastic peak, the sample will be at the focal point of the analyzer (maximum signal and best shaped peak), and all of the Auger peaks will be in the correct positions.